OpenStudy (anonymous):

Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. Provide the solution to the system and explain what the solution represents on the graph. (Equation 1) x + 4y = -16 (Equation 2) 3x + 2y = 12

OpenStudy (fifciol):

convert the equations to the form y=mx +b , then draw these lines in coordinate system. The point (x,y) when they intersect is the solution

OpenStudy (anonymous):

how do I do that ?

OpenStudy (fifciol):

x + 4y =-16 4y = -x -16/:4 y=-1/4x -4 this is our first line. substitute for x some two randomly choosen values and get your y(for example if x=0 y =-4 and if x=4 y=-5) You have then two points (0,-4) and (4,-5). Draw them and then the line that intersects these points. Do that with the other equations. When these line intersects in one point, this point(in this case (8,-6)) is the solution(you read that point from your coordinate system)

OpenStudy (anonymous):

so is (8,-6) the answer ?

OpenStudy (fifciol):

yes :)